OpenStudy (anonymous):

Solve 6 over x minus 4 equals 4 over x for x and determine if the solution is extraneous or not.

OpenStudy (misty1212):

\[\frac{6}{x-4}=\frac{4}{x}\]?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

@misty1212 can you help me?